# Are most sentences unique? An empirical examination of Chomskyan claims

### v1

@@ Abstract
A repeated claim in linguistics is that the majority of linguistic utterances are unique. For example, Pinker (1994: 10), summarizing an argument by Noam Chomsky, states that “virtually every sentence that a person utters or understands is a brand-new combination of words, appearing for the first time in the history of the universe.” With the increased availability of large corpora, this is a claim that can be empirically investigated. The current paper addresses the question by using the NLTK Python library to parse corpora of different genres, providing counts of exact string matches in each. Results show that while completely unique sentences are often the majority of corpora, this is highly constrained by genre, and that duplicate sentences are not an insignificant part of any individual corpus.
@@ 1 Introduction
Counterarguments from “functional” (or “usage based”) theories of language suggest that properties of language arise from general human cognitive processes applied to communication (c.f. Greenberg 1963; Pike 1967; Hawkins 2014). The recent success of large language models in mimicking human language via statistical pattern matching over large datasets of naturalistic texts seem to support such a usage-based approach. However, if the Pinker/Chomsky claim regarding the unique nature of utterances is true, this poses a problem for such usage-based theories.
@@ 2 Methodology
In this investigation the matching criteria was taken to be “exact” string matches. As we are working with written text or written transcriptions of spoken text, two normalization procedures were implemented. The first is to lowercase all words, and the second is to remove punctuation. Since there is no lowercasing or punctuation in spoken communication, the removal of such conventions allows for a more accurate comparison between spoken and written corpora. Resulting normalized strings were then compared.
There are two general measures by which sentences can be compared. The first is within corpora, and the second is across corpora. Each of the corpora measured here contained multiple sub-corpora. That is, the BNC contains multiple documents that represent random samples of text from various sources, while the CHILDES corpora of North American English contains documents from various researchers involving transcriptions of speech from multiple participants. This variation allows for investigation of patterns that may generalize to a particular domain (i.e. within child language acquisition) as well as across genres (i.e. patterns that hold between academic writing and news articles). We can also observe the effect of corpus size on the prevalence of novel utterances.
3 Results and discussion
Table 2 displays the results of an initial analysis of sentences on the basis of exact string matches. In this table we can see that for each individual corpus, the average percentage of sentences that are duplicated within each text varies between 3.8% and 31.5%. When we compare the total number of sentences in a corpus with the total number of duplicates, we find that this range increases from 0.8% to 55.8%.
NLTK corpus Texts Within-text dups (percent) Sentences Within-corpus dups (percent)
brown 500 4.3 57,340 2.2
gutenberg 18 9.5 98,552 6.7
movie_reviews 2,000 12.2 71,532 8.9
webtext 6 8.5 25,733 7.1
inaugural 60 3.8 5,395 0.8
state_union 65 6.1 17,930 3.9
bnc 4,049 7.6 6,026,276 12.5
childes 7,167 31.5 2,202,885 55.8
Table 2: Percentages of duplicate sentences in respective corpora
@@ 3 Results and discussion
The overall low amount of repeated sentences across corpora (excepting the CHILDES dataset) does support the initial claim by Pinker, yet the percentages are not as low as one might suppose from the strength of the claim (“virtually every sentence…”). Given that the CHILDES corpus is the only corpus investigated which is purely transcribed from spoken interactions, it may be that the claim only holds for primarily written text. We might also consider that many interactions are “scripted” (c.f. Frake 1964; Hymes 1986, 1989), such that we would expect to find a higher proportion of duplicated sentences in everyday life. As a final note, the degree of duplication may also be affected by the size of a corpus - higher sentence counts may increase the likelihood of duplicate sentences.

### v3

@@ Abstract
A repeated claim in linguistics is that the majority of linguistic utterances are unique. For example, Pinker (1994: 10), summarizing an argument by Noam Chomsky, states that “virtually every sentence that a person utters or understands is a brand-new combination of words, appearing for the first time in the history of the universe.” With the increased availability of large corpora, this is a claim that can be empirically investigated. The current paper addresses the question by using the NLTK Python library to parse corpora of different genres, providing counts of exact string matches in each. Results show that while completely unique sentences are the majority of corpora, actual percentages are highly constrained by genre, and that duplicate sentences are not an insignificant part of any individual corpus.
@@ 1 Introduction
Counterarguments from “functional” (or “usage based”) theories of language suggest that properties of language arise from general human cognitive processes applied to communication (c.f. Greenberg 1963; Pike 1967; Hawkins 2014). The recent success of large language models in mimicking human language via statistical pattern matching over large datasets of naturalistic texts seem to support such a usage-based approach. However, if the Chomsky/Pinker claim regarding the unique nature of utterances is true, this poses a problem for such usage-based theories.
@@ 2 Methodology
In this investigation the matching criteria was taken to be “exact” string matches. As we are working with written text or written transcriptions of spoken text, two normalization procedures were implemented. The first is to lowercase all words, and the second is to remove punctuation. Since there is no lowercasing or punctuation in spoken communication, the removal of such conventions allows for a more accurate comparison between spoken and written corpora. As a final step, sentences with fewer than three words were excluded. Resulting normalized strings were then compared.
There are two general measures by which sentences can be compared. The first is within corpora, and the second is across corpora. Each of the corpora measured here contained multiple sub-corpora. That is, the BNC contains multiple documents that represent random samples of text from various sources, while the CHILDES corpora of North American English contains documents from various researchers involving transcriptions of speech from multiple participants. This variation allows for investigation of patterns that may generalize to a particular domain (i.e. within child language acquisition) as well as across genres (i.e. patterns that hold between academic writing and news articles). We can also observe the effect of corpus size on the prevalence of novel utterances. [^4] Python code for reproducing these results can be found at: https://github.com/lingdoc/unique_sentences
3 Results
Table 2 displays the results of an initial analysis of sentences on the basis of exact string matches. In this table we can see that for each individual corpus, the average number of duplicates per text is quite low, with the exception of the CHILDES corpus, which reaches nearly 8%. When we compare the total number of sentences in a corpus with the total number of duplicates (across texts), we find that this range increases somewhat, from less than 1% to nearly 30%, depending on the corpus. If we then compute the duplicate percentage based on sentences with three or more words for all sentences across all corpora (nearly 7 million sentences), the number of duplicates rises to roughly 10%. [^5] Thanks to comments and feedback from X/Twitter users on this table and code to produce it, particularly @tlonic.
Corpus Texts Total Sentences Sentences with 3+ words Within-text dups (percent) Across-text/corpus dups (percent)
brown 500 57,340 54,852 0.3 0.58
gutenberg 18 98,552 90,579 1.23 1.17
movie_reviews 2,000 71,532 65,161 0.09 1.82
webtext 6 25,733 23,360 2.96 2.98
inaugural 60 5,395 5,353 0.07 0.19
state_union 65 17,930 17,370 0.19 1.06
bnc 4,049 6,026,276 5,429,986 1.65 5.37
childes 7,167 2,202,885 1,282,614 7.92 28.42
Totals: 13,865 8,505,643 6,969,275 1.8 9.67
Table 2: Sentences and duplicates in respective corpora
4 Discussion
@@
The overall low amount of repeated sentences across corpora (excepting the CHILDES dataset) does seem to support the initial claim by Chomsky/Pinker, yet the percentages are not as low as one might suppose from the strength of the claim (“virtually every sentence…”). In fact, there are duplicate sentences in every single corpus investigated, indicating that repeated sentences are a natural aspect of language in use. Given that the CHILDES corpus is the only corpus investigated which is transcribed solely from spoken interactions, it is possible that the claim holds for primarily written text, though even with purely written text corpora it cannot be maintained in its strongest form.
Work in anthropological linguistics has also shown that many interactions are “scripted” (c.f. Frake 1964; Hymes 1986, 1989), such that we would expect to find a higher proportion of duplicated sentences in everyday life. When ordering a drink at a local shop, for example, there are likely to be a somewhat limited set of interactional/linguistic options for participants. Settings or schemas that are more constrained or often repeated will increase the likelihood of repeated utterances. It also seems that the degree of duplication may also be affected by the size of a corpus - higher sentence counts could increase the likelihood of duplicate sentences. This reflects the nature of “intertextuality” (c.f. Kristeva 1980; Bauman & Briggs 1990; Orr 2003; Still & Worton 1991), whereby texts and interactions often refer to each other.
The findings of this study indicate that the strongest version of the original Chomsky/Pinker claim is false. While the combinatorial possibilities of language do allow for nearly infinite unique utterances, the degree to which a given set of sentences consists of unique strings depends largely on the environment or context in which those sentences were produced. Repetition is in fact a useful device for framing information and guiding attention in interaction (consider dialogue or poetry in scripted performances, for example), and from this perspective it would be surprising if “virtually every” utterance was unique.
There are some caveats that should be made regarding this investigation. I acknowledge, first, that is somewhat limited in scope. We have looked specifically at strings of words/tokens that are three or more words in length, but to what degree can these be considered sentences? There is some argument to be made that the CHILDES corpus is in fact more representative of human language given that it is a spoken corpus of actual interactions, whereas the written corpora are somewhat idealized representations. Additionally, many “sentences” produced in interaction are in fact phrases or fragments (i.e. “constructions) - it could be that breaking sentences into phrases would result in a higher percentage of duplications.
A second point to note is that this investigation focuses only on a single language: English. Although it seems likely that similar principles would hold for other languages, it is possible that these results would change depending on linguistic and cultural context. This raises an interesting thought experiment: is there a language in which investigations of corpora might reveal no duplicate sentences at all? If this is a possibility, what would such a language look like and what would it tell us about human language? While such a question is beyond the scope of this paper, future research could investigate corpora in other languages to explore these ideas further.
As a final point, the low amount of repeated sentences in the corpora explored here highlights the creative nature of language. Combined with the observed increase of repetitions in child-directed speech corpora, it suggests that there is a functional basis for such repetitions. This leads to a theoretical observation that repetition (as referred to by statistical or frequentist models) may be beneficial for language learning, but does not necessarily accurately reflect how language is used more broadly. This mirrors the success of recent Large Language Models, which perform well at recognizable tasks (for which there is robust training data) but not as well on novel or creative tasks (for which there is little training data). As such, it seems we might need a new theory of language that combines elements of the “generative” and “usage-based” approaches (c.f. Rastelli 2025).